OpenStudy (anonymous):

What is the value of x in the equation 0.2(x + 1) + 0.5x = –0.3(x – 4)?

OpenStudy (gorv):

0.2x+0.2+0.5x=-0.3x+1.2 0.7x+0.3x=1.2-0.2 1.0x=1.0 x=1 @nena96

OpenStudy (anonymous):

Thanks